# Testing scalar-tensor-vector gravity theory via black hole photon rings

### v1

@@ III Geodesic Equations and Effective Potentials of Reissner-Nordström Black Holes in Modified Gravity
For a Schwarzschild black hole, Fig. 3 shows $r_{h}=2$2 for the event horizon and $r_{\text{ph}}=3$ for the photon sphere.
@@ IV Fine Structure of the Photon Sphere Orbits
The fine structure of photon-sphere orbits directly reflects the strong-field gravity of a black hole. Near the event horizon, photons propagate along null geodesics and generate a hierarchy of nested sub-rings, whose radii, brightness, and even stability are controlled by intrinsic black-hole parameters such as mass and spin. In this section, we apply backward ray tracing to study photon propagation around the $RN$ black hole in the $STVG$ framework and its interaction with an equatorial thin accretion disk. The disk is modeled as geometrically thin, and the observer is placed along the direction of the black hole’s north pole. Owing to strong gravitational bending, photon trajectories can intersect the disk plane repeatedly, producing intricate, partially overlapping paths. We classify the rays by the number of geodesic–disk intersections and quantify how each class contributes to the total observed intensity. Based on this classification, we compute the cumulative brightness distribution, reconstruct the black-hole image for a background-free emitting disk, and examine how multiple path windings shape the shadow and photon-sphere substructure.
@@
(a)In this figure, the impact parameter $b$ can be interpreted as the radial offset—measured in the observer’s image plane—between the light ray received at infinity and the black-hole center.
@@ V Conclusions and Discussions
This study uncovers the unique physical properties of the photon ring structure of the $RN$ black hole in the $STVG$ theory, and provides an important avenue for the observational verification of quantum gravity theories. The monotonic relationship and non-degeneracy between the $MOG$ parameter $\alpha$, the electric charge $Q$ and the photon sphere radius not only provide an observational criterion for distinguishing different quantum gravity theories (e.g. Saydullayev et al., 2025; Yasmin and Jamil, 2025; Oteev et al., 2025), but also exhibit observable signatures of quantum gravitational effects in the strong gravitational field regime. Parameter constraint results based on $EHT$ observational data show that the theoretical predictions of this theory are in good agreement with existing observations (e.g. Akiyama et al., 2019a, 2022a), yet the current observational precision still limits the accurate discrimination of different values of $\alpha$ and $Q$. Future higher-resolution observational facilities, including next-generation $EHT$ Johnson et al. (2023) and multi-band polarimetric observations Hadar et al. (2022), are expected to enable the precise measurement of the fine structure of photon rings. However, the static, spherically symmetric model and the simplified accretion disk assumption adopted in this study limit the universality of the results to a certain extent. Future work should be extended to Kerr-type black holes with spin effects taken into account, precise magnetohydrodynamic simulations , and plasma effects (e.g. Dihingia and Fendt, 2025; Koide, 2010; Jiménez-Rosales et al., 2021; Kocherlakota et al., 2023; Dhruv et al., 2025). In the context of scalar-tensor-vector gravity theory, the detailed study of the photon ring structure around the Reissner-Nordström black hole paves the way for new experimental approaches to address the fundamental problem of quantum gravity unification. With the continuous progress in observational technologies and the refinement of theoretical models, this research is expected to make significant contributions to the validation of quantum gravity theories, the understanding of black holes’ quantum nature and the exploration of spacetime’s microscopic structure.

### v2

@@ III Geodesic Equations and Effective Potentials of Reissner-Nordström Black Holes in Modified Gravity
For a Schwarzschild black hole, Fig.3 shows $r_{h}=2$ for the event horizon and $r_{\text{ph}}=3$ for the photon sphere.
@@ IV Fine Structure of the Photon Sphere Orbits
The fine structure of photon-sphere orbits directly reflects the strong-field gravity of a black hole. Near the event horizon, photons propagate along null geodesics and generate a hierarchy of nested sub-rings, whose radii, brightness, and even stability are controlled by intrinsic black-hole parameters such as mass and spin. In this section, we apply backward ray tracing to study photon propagation around the $RN$ black hole in the $STVG$ framework and its interaction with an equatorial thin accretion disk. The disk is modeled as geometrically thin, and the observer is placed along the direction of the black hole’s north pole. Owing to strong gravitational bending, photon trajectories can intersect the disk plane repeatedly, producing intricate, partially overlapping paths. We classify the rays by the number of geodesic-disk intersections and quantify how each class contributes to the total observed intensity. Based on this classification, we compute the cumulative brightness distribution, reconstruct the black-hole image for a background-free emitting disk, and examine how multiple path windings shape the shadow and photon-sphere substructure.
@@
(a)In this figure, the impact parameter $b$ can be interpreted as the radial offset-measured in the observer’s image plane-between the light ray received at infinity and the black-hole center.
@@ V Conclusions and Discussions
This study uncovers the unique physical properties of the photon ring structure of the $RN$ black hole in the $STVG$ theory, and provides an important avenue for the observational verification of quantum gravity theories. The monotonic relationship and non-degeneracy between the $MOG$ parameter $\alpha$, the electric charge $Q$ and the photon sphere radius not only provide an observational criterion for distinguishing different quantum gravity theories (e.g. Saydullayev et al., 2025; Yasmin and Jamil, 2025; Oteev et al., 2025), but also exhibit observable signatures of quantum gravitational effects in the strong gravitational field regime. Parameter constraint results based on $EHT$ observational data show that the theoretical predictions of this theory are in good agreement with existing observations (e.g. Akiyama et al., 2019a, 2022a), yet the current observational precision still limits the accurate discrimination of different values of $\alpha$ and $Q$. Future higher-resolution observational facilities, including next-generation $EHT$ Johnson et al. (2023) and multi-band polarimetric observations Hadar et al. (2022), are expected to enable the precise measurement of the fine structure of photon rings. However, the static, spherically symmetric model and the simplified accretion disk assumption adopted in this study limit the universality of the results to a certain extent. Future work should be extended to Kerr-type black holes with spin effects taken into account, precise magnetohydrodynamic simulations , and plasma effects (e.g. Dihingia and Fendt, 2025; Koide, 2010; Jiménez-Rosales et al., 2021; Kocherlakota et al., 2023; Dhruv et al., 2025). In the context of $STVG$ theory, the detailed study of the photon ring structure around the $RN$ black hole paves the way for new experimental approaches to address the fundamental problem of quantum gravity unification. With the continuous progress in observational technologies and the refinement of theoretical models, this research is expected to make significant contributions to the validation of quantum gravity theories, the understanding of black holes’ quantum nature and the exploration of spacetime’s microscopic structure.